OpenStudy (anonymous):

FAN and MEDAL. y = -3.5x - 3.5 y = -3.5x + 3.5 No solution Infinite solutions

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

hold on @sunshinee

OpenStudy (anonymous):

no solution

OpenStudy (anonymous):

since both of the lines are parallel

OpenStudy (anonymous):

to each other* @sunshinee

OpenStudy (anonymous):

@sayakdbz would this be y + 2x = -1 3y - x = 4 -1,1 or 1,1 or parallel no solution

OpenStudy (anonymous):

@unshine\

OpenStudy (anonymous):

@sunshinee

OpenStudy (anonymous):

?

OpenStudy (anonymous):

There would be a unique solution for the question you asked second, it comes out to be x=-1 and y=1. Check for yourself.

OpenStudy (anonymous):

thank you @silversurfer

OpenStudy (anonymous):

44444444444444444444